OpenStudy (ttop0816):

Evaluate |-11| + |-7|. -18 -4 4 18

OpenStudy (anonymous):

or those absolute values?

OpenStudy (anonymous):

11+7 then

OpenStudy (ttop0816):

yes

OpenStudy (anonymous):

absolute values aren't negative so try 11+7, which means takeaway the negatives

OpenStudy (ttop0816):

thank you! do you mind helping me out with another question? @timo86m & @sportcraze0

OpenStudy (anonymous):

Sure I'll see if Ik it

OpenStudy (ttop0816):

Which of the following values is farthest to the left on the number line? 1/5 (0.1)^2 0.02

OpenStudy (ttop0816):

Evaluate a^2 + b for a =1/2 and b = -1/2. -1/2 -1/4 0 1/2

OpenStudy (anonymous):

Ok 1/5 = 0.2, which is greater than 0.02, we r looking for the lowest number so cross out 1/5 (0.1)^2 equals 0.01

OpenStudy (anonymous):

which is greater 0.01or0.02

OpenStudy (ttop0816):

0.02

OpenStudy (ttop0816):

i mean 0.01

OpenStudy (petiteme):

a^2 + b for a =1/2 and b = -1/2. 1/2 ^2 + -1/2 1/4 - 1/2 -1/4

OpenStudy (anonymous):

0.02is greater,but the aner is 0.01

OpenStudy (anonymous):

*answer is 0.01

OpenStudy (candy13106):

do u got the answer yet

OpenStudy (candy13106):

yes im talking to u young lady